The table below sets out the total product plan of a firm operating in a perfectly competitive market. If the firm hires workers at $3 a week. Complete the following table

Labour Output Fixed Cost Variable Cost Total Cost AVC ATC MC
0 0 0 12
1 6
2 15
3 21
4 24
Answer #1
Labor (L) Output (Q) FC ($) VC($)= $3*L TC($)=FC+VC AVC($)=VC/Q ATC($)=TC/Q MC($)=∆TC/∆Q
0 0 12 0 12 0 - -
1 6 12 3 15 0.50 2.50 0.50
2 15 12 6 18 0.40 1.20 0.33
3 21 12 9 21

0.43

1.00 0.50
4 24 12 12 24 0.50 1.00 1.00